Consider two random variables, X and Y. Let E(X) and E(Y) denote the population means of X and Y respectively. Further, let Var(X) and Var(Y) denote the population variances of X and Y. Consider another random variable that is a linear combination of X and Y Z- 3X- Y What is the population variance of Z? Assume that X and Y are independent, which is to say that their covariance is zero.
Answer #1

Given that, population mean and population variance of X and Y are, E(X) , E(Y) and Var(X), Var(Y)

Let Z = 3X - Y

Assume that X and Y are independent, so Cov(X, Y) = 0

We want to find, the Variance of Z,

Var(Z)

= Var ( 3X - Y)

= Var(3X) + Var(Y) - 2 * 3 * Cov(X,Y)

= 32 * Var(X) + Var(Y) - 0

= 9 * Var(X) + Var(Y)

Therefore,

Var(Z) = 9 * Var(X) + Var(Y)